OpenStudy (anonymous):

Graph each function by making a table of coordinates. \[f(x) = 4^x\] I also, have a question...what's a graphing utility?

OpenStudy (aric200):

A graphing Utility is any tool for automatically graphing functions. This could be a graphing calculator, a computer algebra system, a special piece of software just for graphing, or even some sort of mechanical device.

OpenStudy (aric200):

As for graphing your on your own sorry haha... But good luck

OpenStudy (dls):

Hint: f(x) = 4^x passes through (0,1) and (1,4). f(x) = 4^x is an exponential function graph so it will be curved. f(x) = 4^x is an exponential,therefore an always increasing function. Lastly,as x tends to infinite,f(x) tends to infinte.

OpenStudy (anonymous):

Thank you both ♥

OpenStudy (dls):

yw